OpenStudy (anonymous):

the rate of change of y is proportional to y. When x=0, y=6 and when x=4, y=15. Write and solve the differential equation that models this verbal statement. Then find the value of y when x=8.

OpenStudy (nincompoop):

f(0) = 6 and f(4) = 15 correct?

OpenStudy (anonymous):

Yes

OpenStudy (nincompoop):

m = (y2-y1)/(x2-x1) = (15-6)/(4-0) = 9/4 then evaluate at x= 8